OpenStudy (anonymous):

56 is what percent of 800? how do you know when to divide by what?

OpenStudy (squirrels):

Try dividing that 800% to get 100% first

OpenStudy (squirrels):

Then divide that number by 56 and that is your answer.

OpenStudy (imstuck):

Or you could rewrite it in "mathspeak" keeping in mind that "is" means equals, and "of" means to multiply.

OpenStudy (squirrels):

Or you can write it as a proportion

OpenStudy (imstuck):

56=x% * 800-->56=800x 56/800=.07 In percents it is 7%

OpenStudy (squirrels):

\[\frac{ 56 }{ 800 }=\frac{ x }{ 100 }\]

OpenStudy (squirrels):

dont give her the answer bruh, she needs to LURN

OpenStudy (imstuck):

So true...sorry. Got over-excited for a minute...

OpenStudy (squirrels):

all is forgiven brother

OpenStudy (imstuck):

sister...just sayin'

OpenStudy (squirrels):

brother mother sister mister all the same thing right? potato tomato. lol just kidding, my bad, sister :3

OpenStudy (imstuck):

It's all good!

OpenStudy (anonymous):

wait i need to know when how to know what to divide....

OpenStudy (anonymous):

how do i know its not the other way around with different problems?

OpenStudy (anonymous):

im confused

OpenStudy (sasogeek):

okay so... what you need to understand is, when you want to find out what percent of a number A is number B, you do B divided by A times 100. 100 cos we're dealing with percentages. and we're doing B/A cos B is some proportion of A. example. if i say what percentage is 70 out of 100, what i need to know is how much of 70 is in 100, per cent, per 100, that's what it means. so, i do 70 out of 100 = 70/100 = 0.7 0.7 per cent which means per 100, = 0.7 * 100 = 70 so what percentage of 100 is 70? 70%. now if we want to know what percentage of 50 is 25? we'd do 25 out of 50 per cent = (25/50) * 100 =0.5 * 100 =50% do you understand?

OpenStudy (anonymous):

so whatever comes after of is the denominator?

OpenStudy (sasogeek):

the question can be twisted in many ways. you just have to understand the concept. for now, you can assume what you just said to be always true. but do you get the whole idea?

OpenStudy (anonymous):

yea im just confused of where what goes when it comes to the is and of

OpenStudy (sasogeek):

okay, try solving your question. use my guidelines, let me know what answer you get

OpenStudy (anonymous):

7%

OpenStudy (sasogeek):

correct... but try doing this... 5 oranges and 10 apples in a fruit basket. what's the percentage of apples in the fruit basket?

OpenStudy (anonymous):

66.6%

OpenStudy (sasogeek):

you're good to go xD